An individual purchases a $82,000 term life insurance with an annual payment of $620. The probability that the individual will survive the year is 0.999157. Find the expected value of the policy for the insurance company. Show your work to receive credit
Answer #1

let X be the net profit for the insurance company

P(X =620) =P(survive) =0.999157

P(X=620-82000=-81380)=P(not survive) =1-0.999157 =0.000843

therefore expected value of the policy for the insurance company E(x)=xP(x)

=620*0.999157-81380*0.000843

=550.87